OpenStudy (anonymous):

a fisherman hooks a trout and reels in his line at 4in/s. Assume the tip of the fishing rod is 12ft above the water and the fish is pulled horizontally directly towards the fisherman. Find the horizontal speed of the fish when it is 20 ft from the fisherman.

OpenStudy (anonymous):



OpenStudy (anonymous):

differentiate L with respect to time to find dL/dt solve for dx/dt plug in the values you are given to find dx/dt at a specific time ( when dL/dt = 1/3 and x= 20

OpenStudy (anonymous):

make sense?

OpenStudy (anonymous):

how did you get dL/dt= 1/3?

OpenStudy (anonymous):

converted to ft/sec